# MetaLLM: A High-performant and Cost-efficient Dynamic Framework for Wrapping LLMs

### v2

@@ Abstract
The rapid progress in machine learning (ML) has brought forth many large language models (LLMs) that excel in various tasks and areas. These LLMs come with different abilities and costs in terms of computation or pricing. Since the demand for each query can vary, e.g., because of the queried domain or its complexity, defaulting to one LLM in an application is not usually the best choice, whether it is the biggest, priciest, or even the one with the best average test performance. Consequently, picking the right LLM that is both accurate and cost-effective for an application remains a challenge. In this paper, we introduce MetaLLM, a framework that dynamically and intelligently routes each query to the optimal LLM (among several available LLMs) for classification tasks, achieving significantly improved accuracy and cost-effectiveness. By framing the selection problem as a multi-armed bandit, MetaLLM balances prediction accuracy and cost efficiency under uncertainty. Our experiments, conducted on popular LLM platforms such as OpenAI’s GPT models, Amazon’s Titan, Anthropic’s Claude, and Meta’s LLaMa, showcase MetaLLM’s efficacy in real-world scenarios, laying the groundwork for future extensions beyond classification tasks [^1] Our code will be published soon at https://github.com/mail-research/MetaLLM-wrapper/..
@@ 1 Introduction
Large language models have shown extraordinary zero-shot capabilities in various tasks and domains, such as text classification, summarization, question answering, and chatbot [22, 1, 25, 32, 21, 29]. Recent works [13, 5, 10] suggest exhaustively scaling the model size and training data size to improve the performance of language models and provoke their emergent abilities; for example, GPT-4, with over $1.74$ trillion parameters, achieves superior performance in several tasks but also incur high economic costs. While this trend of scaling language models will continue in the future, there is also a growing diversification in recent models in terms of task or (sub-)domain specialization and computational costs. Consequently, for model users, identifying which LLM is best suited for their applications has become a daunting task. When factoring in the cost constraints, either computational resources or API service pricing, this task becomes even more complex.
@@
In this paper, we imagine a world with several LLM providers, such as OpenAI [^2] https://platform.openai.com/docs/models or Amazon’s Bedrock [^3] https://aws.amazon.com/bedrock; each provides service access to a diverse suite of LLMs with heterogeneous capabilities and cost structures. Here, an LLM user asks an important question: How do I select an LLM $i$ (out of $k$ LLMs) for optimal performance and usage cost in my application? One option is combining multiple LLMs as seen in existing ensemble methods [12, 30], but this approach will yield significantly higher service costs; another option is cascading over a set of LLMs [4], but such an approach still requires querying the LLMs until we could find the best one. On the other hand, defaulting to a single LLM to avoid the extra cost of querying multiple models – by predicting the performance of LLMs [27, 17] – may not also be optimal; for some queries, a less expensive model may also provide correct answers. Furthermore, as different LLMs exhibit very distinctive abilities on different tasks and data distribution [12, 30], for a query, it is possible for LLM $i$ to perform better than LLM $j$ even though the cost of LLM $i$ is noticeably lower than the cost of $j$.
@@
In this work, we introduce a framework that wraps around a set of LLMs with diverse capabilities and brings the best performance at a more affordable cost to a user’s application. The underlying component of this framework is a multi-arm bandit algorithm that routes each query to the least expensive LLM with the correct answer for the zero-shot classification task. Essentially, we formulate the problem of selecting the “best” LLM for a query as a decision-making problem under uncertainty: when a new query arrives, the proposed framework picks an LLM to answer and observe whether the LLM provides a correct answer; the goal is to maximize the total reward – a specific trade-off between the overall performance and usage cost. Intuitively, our method prefers the LLM that has a high probability of answering a query at a low cost. Our contributions can be summarized as follows:
@@
  • Our experimental results on benchmark datasets and popular API services, including OpenAI and Amazon’s Bedrock, demonstrate the ability of MetaLLM in identifying the optimal LLM in terms of cost and performance. Specifically, MetaLLM improves the accuracy of the best model by around $1\%$ while saving up to $60\%$ and $40\%$ of the total price on OpenAI and Bedrock APIs, respectively.
The zero-shot text classification is an important task in many NLP applications, and LLMs are increasingly used as the predictive models for this task. Employing LLMs removes the users from several challenges when creating a text classification model from scratch, including imbalanced data, missing data, corrupted data, or zero-shot capability. While our work focuses on zero-shot classification problems, the MetaLLM framework can be extended to arbitrary language tasks by modifying the reward function to incorporate suitable metrics assessing the quality of the responses. However, we leave it for future work.
@@ 2 Related Works
Large Language Models. The emergence of large language models (LLMs) has fundamentally transformed several domains, including natural language processing, computer vision, and e-commerce, and on diverse tasks such as (zero-shot) classification, question-answering, and recommendation [20, 15, 16]. The impressive effectiveness and versatility of these LLMs have come at the price of a drastic increase in LLM sizes, along with significant computational costs and data required to train, and expensive computational resources to perform inference with them. Consequently, several companies or services are now offering users access to LLMs with diverse capabilities and heterogeneous cost structures. For instance, the costs for processing 10 million tokens are \$15, \$30, \$75, and \$80 for Titan Lite, Cohere Command-Light, Llama 2 7B, and Claude Instant, respectively, using Amazon Bedrock APIs. This abundance of API choices significantly burdens their users with the decision “which LLM should I use?”, as different LLM APIs are known for their diverse capabilities for the prediction tasks [14, 18, 19].
@@
Prompt Optimization and Mixture of Experts. Fine-tuning is a standard option to improve the performance of LLMs for specific tasks. Mixture-of-Experts (MoE) [8, 26, 7, 28] trains a routing operator within the large model to enhance its performance; essentially, MoE assumes the model as a collection of “experts” (modules) and learn to route the input to the best expert. These approaches require training the LLMs, which is challenging for many users, while their single-LLM enhancements are usually model and scenario specific. Prompting reasoning like Chain-of-Thought  [32, 31, 34] or Tree of Thoughts [33] could improve the LLM performance without additional training. Both MoE and prompt-based reasoning, however, could not benefit from the large number of available LLMs, some of which could be significantly less expensive to use.
@@
Model Ensemble. LLMs have been shown to yield diverse capabilities due to their architecture and dataset [12, 30]. Jiang et al. [12] observe that over $5000$ instructions, the optimal LLM for each query significantly varies, and there is no single optimal model regardless of their size. Therefore they recommend ensembling to combine the strength of multiple LLMs for a better performance. FrugalML [2, 3] cascades multiple machine learning models by querying sequentially until getting a response with a high confidence score. Motivated by FrugalML, Chen et al. [4] exploit the full potential of LLMs by applying multiple techniques, including prompt engineering, caching, and cascading, for higher-quality answers.
@@
Model Selection. Different from previous approaches that combine the strength of multiple models, there are some attempts to query a single LLM for each task. Hari and Thomson [9] suggest training a language model to predict the performance of LLMs and route to the model with the highest performance, which incorporates the high cost of the router. Šakota et al. [24] formulate the cost-performance trade-off as an integer linear programming (ILP) problem and use existing ILP solvers to assign appropriate LLMs for each input query. Lu et al. [17] distill the preference of off-the-shelf reward models to select the LLM that achieves the best performance, ignoring its cost. Ding et al. [6] train a router that maps easy queries to small models and hard queries to large models; however, their framework is only applicable when we have two LLMs. Recently, Hu et al. [11] proposed a benchmark to evaluate routing models by their cost and performance on downstream tasks.
@@
Commercialized Foundation Model Services. Recently, many companies have commercialized their LLM models so that customers can apply or even fine-tune LLMs for their own use cases without heavy technical knowledge. In 2023, OpenAI released many language models, such as text-ada-001, text-babbage-001, text-curie-001, and text-davinci-002444These models have been deprecated since 2024-01-04., with various prices and capabilities. Their users can use these models as a chatbot, extract text embedding, or even ask them to write code. Recently, Amazon also brought forth their Bedrock APIs that allow access to several foundation generative models, including text generation, image generation, and multimodal models. Bedrock provides Amazon’s Titan models and also models from other companies, such as Anthropic, Cohere, Meta, Mistral, and Stability.
@@ 3.1 Preliminaries
LLM APIs. We consider the case where the user has access to a set $K$ of $k$ different LLM APIs; each LLM $M_{i}$ has a different capability, determined by how well $M_{i}$ can answer a query $x$, and a cost $c_{i}$. For zero-shot classification, we represent the capability $a_{i}(x)\in\{0,1\}$ of an LLM $M_{i}$ on a sample $x$ by comparing the answer to its corresponding ground-truth label: $a_{i}(x)=1$ if $M_{i}(x)=y$. Usually, the model with a higher cost has better capability. If the user has a limited budget, they can choose less expensive models for their application. In contrast, if they require better performance, more expensive models will be more desirable. However, in general, the more expensive model is not always the best-performing choice for all queries [12, 30]
Problem Formulation. The goal of MetaLLM  is to learn a routing function $f:\mathcal{X}\to K$ that dispatches a query $x$ to an appropriate LLM to achieve a good response with a lower cost. For example, given a subset of LLMs, $K^{\prime}\subseteq K$, that can give good responses for $x$, MetaLLM’s objective is to return $\arg\min_{l\in K^{\prime}}c_{l}$. In practice, the user wants the ability to balance the performance and the usage cost, depending on the needs of their application. More specifically, given a budget $b$, the user wants to maximize the performance while spending less than $b$ to query the APIs.
@@ 3.2 The optimal solution of routing objective
Let $S\in\ \{0,1\}^{N\times k},A\in\mathbb{R}^{N\times k}$, where $S$ is a matrix that has $S_{i,j}=1$ if and only if $f(x_{i})=j$, and $A_{i,j}=a_{j}(x_{i})$ is the performance of the $j$-th API on the sample $x_{i}$. The objective in 1 can be reformulated as follows:
@@
We can relax the constraint of $S$ to $S\in\mathbb{R}^{N\times k}$ and $\sum_{j=1}^{k}S_{i,j}=1$, and its dual problem can be derived as
@@
Chen et al. [3] study similar formulation for model cascading and suggest solving the dual form 3. In the case where we have an exact accuracy matrix $A$, we can optimize 2 to find the optimal $S$ and choose the suitable API $j$ such that $S_{i,j}=1$.
@@ 3.3 The Proposed MetaLLM
When deploying an application, we cannot know the exact accuracy of an LLM’s API on a test sample before sending the test sample through this LLM. Previous works [3, 2] learn a model to predict the performance of each API. These methods cascade multiple machine learning models and query them iteratively until the response has high confidence; consequently, they are highly expensive, especially when there are a substantial number of queries.
In this work, we approach this problem from a different perspective. Instead of training an accuracy predictor, we reformulate this problem as a multi-armed bandit. Specifically, for each input query, we define an LLM as an “arm”, and obtain a reward expressing the performance of the LLM and the cost for that query. The benefit of this formulation is twofold: first, it allows the modeler to focus on designing the reward function to capture their application needs, making the framework more versatile; second, we can take advantage of the extensive and well-developed research on multi-arm bandit, including their rigorous theoretical foundations and practical solutions. The MetaLLM framework is depicted in Figure 1.
The remaining question will discuss our design of the reward function for the zero-shot classification task. Chen et al. [3] prove that if a cost scaling $p\in\mathbb{R}$ is the solution of 3, the routing function $f(x)=\arg\max_{i}a_{i}(x)-pc_{i}$ will be the optimal solution of 1. Intuitively, this strategy prefers the LLM with high performance and low-cost value. Motivated by that theoretical result, we propose the following reward function for training the multi-armed bandit:
@@
During training, we compute the expected reward $Q_{j^{\prime}}(x;\theta)$ of each arm $j^{\prime}$ for a training sample $x$ and choose the arm $j$ with the highest expected reward and minimize the objective
$\theta=\arg\min_{\theta^{\prime}}\|Q_{j}(x_{i},\theta^{\prime})-(a_{j}(x_{i})-p(c_{j}))\|_{2}^{2}.$ (5)
During inference, MetaLLM returns the expected reward of each arm for a query and dispatches that query to the LLM with the highest reward. The training and inference algorithm for MetaLLM is illustrated in Algorithm 1.
Algorithm 1 Our MetaLLM framework
input k LLMs, each has cost $c_{j}$ and accuracy function $a_{j}(\cdot)$, cost scaling $p$, parameters $\theta$, training sample $\{x_{1},\dots,x_{N}\}$, test query $x$.
output The optimal LLM for the query $x$.
### Training step
for s in 1..step do
     for i in 1..N do
         $j\leftarrow\arg\max_{j}^{\prime}Q_{j^{\prime}}(x_{i};\theta)$
         $\theta\leftarrow\arg\min_{\theta^{\prime}}\|Q_{j}(x_{i},\theta^{\prime})-(a_{j}(x_{i})-pc_{j})\|_{2}^{2}$
     end for
end for
### Inference step return $\arg\max_{j}Q_{j}(x;\theta)$

### v3

@@ Abstract
The rapid progress in machine learning (ML) has brought forth many large language models (LLMs) that excel in various tasks and areas. These LLMs come with different abilities and costs in terms of computation or pricing. Since the demand for each query can vary, e.g., because of the queried domain or its complexity, defaulting to one LLM in an application is not usually the best choice, whether it is the biggest, priciest, or even the one with the best average test performance. Consequently, picking the right LLM that is both accurate and cost-effective for an application is necessary yet remains a challenge. In this paper, we introduce MetaLLM, a framework that dynamically and intelligently routes each query to the optimal LLM (among several available LLMs) for classification and multi-choice question-answering tasks, achieving significantly improved accuracy and cost-effectiveness. By framing the selection problem as a multi-armed bandit, MetaLLM balances prediction accuracy and cost efficiency under uncertainty. Our experiments, conducted on popular LLM platforms such as OpenAI and Together AI, as well as open-source LLM, showcase MetaLLM’s efficacy in real-world scenarios, laying the groundwork for future extensions [^1] Our code is available here..
@@ 1 Introduction
Large language models have shown extraordinary zero-shot capabilities in various tasks and domains, such as text classification, summarization, question answering, and chatbot (Radford et al., 2019; Brown et al., 2020; Schick & Schütze, 2021; Wei et al., 2022; Ouyang et al., 2022; Trivedi et al., 2023). Recent works (Kaplan et al., 2020; Chowdhery et al., 2023; Hoffmann et al., 2022) suggest exhaustively scaling the model size and training data size to improve the performance of language models and provoke their emergent abilities; for example, GPT-4, with over $1.74$ trillion parameters, achieves superior performance in several tasks but also incurs high economic costs. While this trend of scaling language models will continue in the future, there is also a growing diversification in recent models in terms of task or (sub-)domain specialization and computational costs. As a consequence, no single model –even the largest and most expensive – can consistently yield the best performance across all tasks. This means that, for model users, identifying which LLM is best suited for their applications will become crucial. However, this is a challenging task, especially when we factor in cost constraints, either in terms of computational resources or API service pricing.
@@
We imagine a world with several LLM providers, such as OpenAI [^2] https://platform.openai.com/docs/models or Together AI [^3] https://www.together.ai/; each provides service access to a diverse suite of LLMs with heterogeneous capabilities and cost structures. Here, an LLM user asks this important question: How do I select an LLM $i$ (out of $k$ LLMs) for optimal performance and usage cost in my application? One option is combining multiple LLMs as seen in existing ensemble methods (Jiang et al., 2023; Wang et al., 2023; Ong et al., ), but this approach will yield significantly higher service costs; another option is cascading over a set of LLMs (Chen et al., 2023), but such an approach still requires querying the LLMs until we could find the best one. On the other hand, defaulting to a single LLM to avoid the extra cost of querying multiple models – by predicting the performance of LLMs (Shnitzer et al., 2023; Lu et al., 2023) – may not also be optimal; for some queries, a less expensive model may also provide correct answers. Furthermore, as different LLMs exhibit very distinctive abilities on different tasks and data distribution (Jiang et al., 2023; Wang et al., 2023), for a query, it is possible for LLM $i$ to perform better than LLM $j$ even though the cost of LLM $i$ is noticeably lower than the cost of $j$.
@@
In this work, we introduce a framework that wraps around a set of LLMs with diverse capabilities and brings the best performance at a more affordable cost to a user’s application. The underlying component of this framework is a multi-arm bandit algorithm that routes each query to the least expensive LLM with the correct answer for the zero-shot classification and question answering task. Essentially, we formulate the problem of selecting the “best” LLM for a query as a decision-making problem under uncertainty: when a new query arrives, the proposed framework picks an LLM to answer and observe whether the LLM provides a correct answer; the goal is to maximize the total reward – a specific trade-off between the overall performance and usage cost. Intuitively, our method prefers the LLM that has a high probability of answering a query at a low cost. Our contributions can be summarized as follows:
@@
  • Our experimental results on benchmark datasets and popular API services, including OpenAI and Together AI, and both closed-source and open-source LLMs, demonstrate the ability of MetaLLM in identifying the optimal LLM in terms of cost and performance. Specifically, MetaLLM improves the accuracy of the best model by around $1\%$ while saving up to $60\%$ and $10\%$ of the total price on OpenAI and Together AI APIs, respectively.
Our work focuses on zero-shot classification and multiple-choice question-answering problems, which are common evaluation choices in the related work (Ong et al., ). Nevertheless, the MetaLLM framework can be extended to arbitrary language tasks by modifying the reward function to incorporate suitable metrics assessing the quality of the responses. We leave it for future work.
@@ 2 Related Works
Large Language Models. The emergence of large language models (LLMs) has fundamentally transformed several domains, including natural language processing, computer vision, and e-commerce, and diverse tasks such as (zero-shot) classification, question-answering, and recommendation (Menghani, 2023; Liu et al., 2023a; b). The impressive effectiveness and versatility of these LLMs have come at the price of a drastic increase in LLM sizes, along with significant computational costs and data required to train, and expensive computational resources to perform inference with them. Consequently, several companies or services are now offering users access to LLMs with diverse capabilities and heterogeneous cost structures. For instance, the costs for processing 10 million tokens are \$0.1, \$0.18, \$0.2, and \$0.3 for Gemma, Llama, Mistral, and Qwen, respectively, using Together AI APIs. This abundance of API choices significantly burdens their users with the decision “which LLM should I use?”, as different LLM APIs are known for their diverse capabilities for the prediction tasks (Liang et al., 2023; McKenzie et al., ; McKenzie et al., 2022).
@@
Prompt Optimization and Mixture of Experts. Fine-tuning is a standard option to improve the performance of LLMs for specific tasks. Mixture-of-Experts (MoE) (Eigen et al., 2013; Shazeer et al., 2017; Du et al., 2022; Si et al., 2023) trains a routing operator within the large model to enhance its performance; essentially, MoE assumes the model as a collection of “experts” (modules) and learn to route the input to the best expert. These approaches require training the LLMs, which is challenging for many users, while their single-LLM enhancements are usually model and scenario specific. Prompting reasoning like Chain-of-Thought  (Wei et al., 2022; Wang et al., 2022; Zhou et al., 2022) or Tree of Thoughts (Yao et al., 2024) could improve the LLM performance without additional training. Both MoE and prompt-based reasoning, however, could not benefit from the large number of available LLMs, some of which could be significantly less expensive to use.
@@
Model Ensemble. LLMs have been shown to yield diverse capabilities due to their architecture and dataset (Jiang et al., 2023; Wang et al., 2023). Jiang et al. (2023) observe that over $5000$ instructions, the optimal LLM for each query significantly varies, and there is no single optimal model regardless of their size. Therefore they recommend ensembling to combine the strength of multiple LLMs for a better performance. FrugalML (Chen et al., 2020; 2022) cascades multiple machine learning models by querying sequentially until getting a response with a high confidence score. Motivated by FrugalML, Chen et al. (2023) exploit the full potential of LLMs by applying multiple techniques, including prompt engineering, caching, and cascading, for higher-quality answers.
@@
Model Selection. Different from previous approaches that combine the strength of multiple models, there are some attempts to query a single LLM for each task. Hari & Thomson (2023) suggest training a language model to predict the performance of LLMs and route to the model with the highest performance, which incorporates the high cost of the router. Šakota et al. (2024) formulate the cost-performance trade-off as an integer linear programming (ILP) problem and use existing ILP solvers to assign appropriate LLMs for each input query. Lu et al. (2023) distill the preference of off-the-shelf reward models to select the LLM that achieves the best performance, ignoring its cost. Ding et al. (2024) train a router that maps easy queries to small models and hard queries to large models; however, their framework is only applicable when we have two LLMs. Recently, Hu et al. (2024) proposed a benchmark to evaluate routing models by their cost and performance on downstream tasks.
@@
Commercialized Foundation Model Services. Recently, many companies have commercialized their LLM models so that customers can apply or even fine-tune LLMs for their own use cases without heavy technical knowledge. In 2023, OpenAI released many language models, such as text-ada-001, text-babbage-001, text-curie-001, and text-davinci-002444These models have been deprecated since 2024-01-04., with various prices and capabilities. Their users can use these models as a chatbot, extract text embedding, or even ask them to write code. Recently, Together AI also brought forth their APIs that allow access to several foundation generative models, including text generation, image generation, and multimodal models, such as Llama, Gemma, Mistral, Qwen, Flux, etc.
@@ 3.1 Preliminaries
LLM APIs. We consider the case where the user has access to a set $K$ of $k$ different LLM APIs; each LLM $M_{i}$ has a different capability, determined by how well $M_{i}$ can answer a query $x$, and a cost $c_{i}$. For zero-shot classification, we represent the capability $a_{i}(x)\in\{0,1\}$ of an LLM $M_{i}$ on a sample $x$ by comparing the answer to its corresponding ground-truth label: $a_{i}(x)=1$ if $M_{i}(x)=y$. Usually, the model with a higher cost has better capability. If the user has a limited budget, they can choose less expensive models for their application. In contrast, if they require better performance, more expensive models will be more desirable. However, in general, the more expensive model is not always the best-performing choice for all queries (Jiang et al., 2023; Wang et al., 2023), making the LLM selection problem significantly challenging.
Problem Formulation. The goal of MetaLLM  is to learn a routing function $f:\mathcal{X}\to K$ that dispatches a query $x$ to an appropriate LLM to achieve a good response with a lower cost. For example, given a subset of LLMs, $K^{\prime}\subseteq K$, that can give good responses for $x$, MetaLLM’s objective is to return $\arg\min_{l\in K^{\prime}}c_{l}$. In practice, the user wants the ability to balance the performance and the usage cost, depending on the needs of their application. More specifically, given a budget $b$, the user wants to maximize the performance on $N$ queries while spending less than $b$ to query the APIs.
@@ 3.2 The optimal solution of routing objective
Let $S\in\{0,1\}^{N\times k}$, $S_{i,j}=\mathbf{1}[f(x_{i})=j]$, represent the choice of the routing function $f$ on $N$ queries, and $A\in\mathbb{R}^{N\times k}$, $A_{i,j}=a_{j}(x_{i})$, represent the performance of the $j$-th API on each sample $x_{i}$. Eqn. (1) aims to choose the highest performance LLM within a cost constraint $b$ and can be reformulated as follows:
@@
Instead of solving a discrete optimization problem, we relax $S\in\{0,1\}^{N\times k}$ to $S\in\mathbb{R}^{N\times k}$, $\sum_{j=1}^{k}S_{i,j}=1$, and has the following dual problem (Boyd & Vandenberghe, 2004):
@@
Chen et al. (2022) study a similar formulation for model cascading and suggest solving the dual form as in Eqn. (3). In the case where we have an exact accuracy matrix $A$, we can optimize Eqn. (2) to find the optimal $S$ and choose the suitable API $j$ such that $S_{i,j}=1$.
@@ 3.3 The Proposed MetaLLM
Algorithm 1 MetaLLM framework
input: $k$ LLMs, each with cost $c_{j}$ and accuracy $a_{j}(\cdot)$; cost scaling $p$, parameters $\theta$, training sample $\{x_{1},\dots,x_{n}\}$, test query $x$.
output: The optimal LLM for the query $x$.
### Initialize the policy
for j in 1..k do
     $A_{j}\leftarrow(X^{\intercal}X+\lambda I)$, $b_{j}\leftarrow X^{\intercal}y_{j}$
     $\theta^{0}_{j}\leftarrow A_{j}^{-1}b_{j}$
end for
### Inference
$j\leftarrow\arg\max_{j^{\prime}}x^{\intercal}\theta_{j^{\prime}}+\alpha\sqrt{x^{\intercal}A_{j^{\prime}}^{-1}x}$
$A_{j}\leftarrow A_{j}+xx^{\intercal}$, $b_{j}\leftarrow b_{j}+r(x,j)x$
$\theta_{j}\leftarrow A_{j}^{-1}b_{j}$ return j
When deploying an application, we cannot know the exact accuracy of an LLM’s API on a test sample before sending the test sample through this LLM. Previous works (Chen et al., 2022; 2020) learn a model to predict the performance of each API. These methods cascade multiple machine learning models and query them iteratively until the response has high confidence; consequently, they are highly expensive, especially when there are a substantial number of queries.
In this work, we approach this problem from a different perspective. Instead of training an accuracy predictor, we reformulate this problem as a multi-armed bandit. Specifically, for each input query, we define an LLM as an “arm”, and obtain a reward expressing the performance of the LLM and the cost for that query. The benefit of this formulation is twofold: first, it allows the modeler to focus on designing the reward function to capture their application needs, making the framework more versatile; second, we can take advantage of the extensive and well-developed research on multi-arm bandits, including their rigorous theoretical foundations and practical solutions. The MetaLLM framework is depicted in Figure 1.
Reward function. The remaining question will discuss our design of the reward function that takes the input query and returns the reward of choosing the $i$-LLM. Chen et al. (2022) prove that if a cost scaling $p\in\mathbb{R}$ is the solution of Eqn. (3), the routing function $f(x)=\arg\max_{i}a_{i}(x)-pc_{i}$ will be the optimal solution of Eqn. (1). Intuitively, this strategy prefers the LLM with high performance and low-cost value. Motivated by that theoretical result, we propose the following reward function for training the multi-armed bandit:
@@
Multi-arm bandit. We assume that the user has access to a training dataset of $n$ queries and can obtain the training performance of each LLM on this dataset. We initialize the reward model $Q_{j}(x;\theta)$ of each arm $j$ by minimizing the objective $\theta_{j}=\arg\min_{\theta}\|Q_{j}(x_{i},\theta)-(a_{j}(x_{i})-pc_{j})\|_{2}^{2}.$
Similar to (Li et al., 2010), we employ a linear reward model $Q_{j}(x;\theta_{j})=x^{\intercal}\theta_{j}$ and obtain the closed-form solution of ridge regression $\theta^{0}_{j}=\big{(}{A^{0}_{j}}\big{)}^{-1}b^{0}_{j},A^{0}_{j}=(X^{\intercal}X+\lambda I),b^{0}_{j}=X^{\intercal}y_{j}$, where $X$ is the data matrix and $y_{j,i}=a_{j}(x_{i})-pc_{j}$.
Given a new query $x_{t}$, we choose an arm using the UCB selection strategy:
$j=\arg\max_{j^{\prime}}x_{t}^{\intercal}\theta^{t}_{j^{\prime}}+\alpha\sqrt{x_{t}^{\intercal}{A^{t}_{j^{\prime}}}^{-1}x_{t}},\quad\alpha>0.$
and update the reward function of the chosen arm, as follows:
$\theta^{t+1}_{j}=\big{(}{A^{t+1}_{j}}\big{)}^{-1}b^{t+1}_{j},A^{t+1}_{j}=A^{t}_{j}+x_{t}x_{t}^{\intercal},b^{t+1}_{j}=b^{t}_{j}+r(x_{t},j)x_{t}.$